OpenStudy (anonymous):

A cylinder is full at 471 cm3 and has a radius of 5 cm. It currently contains 314 cm3 of water. What will the difference in the height of the water be when the cylinder is full? Use 3.14 to approximate pi.

OpenStudy (campbell_st):

well use the formula for volume of a cylinder \[V = \pi r^2 h\] making h the subject of the equation its \[h = \frac{V}{\pi r^2}\] you know r = 25 and V = 471 and also V = 314 so substitute the information to find the height of the full container and the other height. Last task is to find the difference in heights

OpenStudy (anonymous):

r=5cm!!!!